A stock you are evaluating just paid an annual dividend of $2.70. Dividends have grown at a constant rate of 2.4 percent over the last 15 years and you expect this to continue.

a. If the required rate of return on the stock is 12.8 percent, what is its fair present value?
b. If the required rate of return on the stock is 15.8 percent, what should the fair value be four years from today?
  
(For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16))
Answer #1

a.Present value=D1/(Required return-Growth rate)

=(2.7*1.024)/(0.128-0.024)

=$26.58(Approx).

b.Present value=D1/(Required return-Growth rate)

=(2.7*1.024)/(0.158-0.024)

=$20.63283582

Hence P4=Present value*(1+Growth Rate)^4

=$20.63283582*(1.024)^4

=$22.69(Approx).